# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ Abstract
A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communications (ISAC) system is proposed to enhance multi-target sensing and multi-user communication. Unlike the conventional RIS, the fRIS incorporates movable elements whose positions can be flexibly adjusted to provide extra spatial degrees of freedom. In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and communication symbol estimation error by optimizing the symbol estimator, transmit beamformer, fRIS phase shifts, and element positions. To solve this problem, an algorithm based on alternating minimization is devised, where subproblems are solved leveraging augmented Lagrangian method, quadratic programming, semidefinite-relaxation, and majorization-minimization techniques. A key challenge exists that the fRIS element positions affect both the incident and reflective channels, leading to the high-order composite functions regarding the positions. As a remedy, it is proved that the high-order terms can be transformed to linear and linear-difference forms using the characteristics of fRIS and structural channels, which facilitates the position optimization. Numerical results validate the effectiveness of the proposed scheme as compared to the conventional RIS-aided ISAC systems and other benchmarks.
Index Terms:
Alternating minimization, fluid reconfigurable intelligent surface, integrated sensing and communications, mismatch and estimation error
@@ I Introduction
In the era of sixth-generation (6G) wireless communications, integrated sensing and communications (ISAC) emerged as a promising paradigm [1], addressing spectrum scarcity and reducing resource consumption. The ISAC aims to incorporate the sensing systems and the communication systems at different levels. Herein, the two systems shared the same spectrum and coexisted with minimal mutual interference [2] using null space projection [3], transceiver design [4], etc. To further mitigate interference, collaborations between the two systems were achieved by sharing critical information    [5], [6]. Additionally, considering the similarities in hardware and signal processing, the recent ISAC advancement focuses on designing unified platforms that can perform the dual functions [7]. For example, in [8], the transmit beamformer was designed by combining weighted radar waveforms and communication symbols. Besides, a wideband scenario about joint design was considered in [9], where mutual information was utilized as a unified metric to design the ISAC waveform.
Meanwhile, reconfigurable intelligent surface (RIS) garnered great attention in the 6G innovation. Specifically, the RIS consists of numerous reflective elements, each of which is able to induce phase shifts to the incident signals. This enables the RIS to dynamically adapt the channel propagation environment by designing the phase shifts [10]. Leveraging this characteristic, work [11] demonstrated significant improvement in communication sum-rate in RIS-aided multi-user scenarios. Apart from this, it was also shown that the RIS can suppress the undesired signals, such as eavesdropping  [12] and jamming [13]. Additionally, the RIS was applied in other diverse scenarios, including radar detection [14], near-field communications [15], and wireless power transfer and communications [16].
To harness the combined benefits of ISAC and RIS, extensive research was conducted on RIS-aided ISAC systems [17]. In [18], authors investigated an echo signal-to-noise ratio (SNR) maximization problem in a RIS-aided ISAC system, which serves a communication user and tracks a target through joint design of the transmit precoder and RIS beamformer. Building on this, [19] extended this framework to multi-user and multi-target scenarios under practical constraints. Furthermore, authors in [20] and [21] considered the energy efficiency maximization for the RIS-aided ISAC systems without compromising the communication and sensing performance. Additionally, researchers examined various other applications of the RIS-aided ISAC systems, such as simultaneously transmitting and reflecting surface [22], unsupervised learning [23], and extended reality [24].
@@
While the RIS-aided ISAC systems offer significant enhancements for 6G communications, there remains room for the improvement in spatial degrees of freedom (DoFs) of the systems and the reduction in system complexity. Particularly, the antennas of the conventional arrays are fixed after assembly, which restricts the spatial diversity [25]. On the other hand, to achieve excellent performance, a massive number of elements are required, which increases manufacturing costs and exacerbates high-dimensional optimization complexity. To address these limitations, movable antenna, also known as fluid antenna, becomes a promising candidate. Enabled by the advancement of software-controllable fluidic metamaterials, the movable antenna can dynamically adjust its position in real time, offering additional spatial DoFs  [26], [27]. Moreover, the enhanced performance brought by the movable antenna allows comparable gains with fewer elements, significantly reducing the computational complexity caused by high-dimensional optimization.
The potential of the fluid antenna system was extensively studied. For instance, researchers demonstrated that it can promote the capacity improvement in point-to-point communication systems [25] and enhance downlink sum-rate in multi-user communication scenarios [28]. Besides, it was applied in ISAC systems to jointly optimize transceiver designs and antenna positions, enhancing communication and sensing performance [29], [30], [31]. Meanwhile, efforts to integrate the RIS with the fluid antenna system was made, where authors in [32] and [33] analyzed the outage probability bounds and asymptotic distributions for fluid antenna-aided RIS systems. Nevertheless, these works still rely on fixed RIS elements and do not fully exploit the spatial DoFs of RIS elements. By contrast, [34] introduced movable RIS elements with flexible positioning and studied the elimination of phase distribution offset in the system. However, this work was limited to a simple single-antenna and single-user communication setup, while did not consider the system optimization problem. This leaves room for further exploration in the potential of the movable RIS elements in ISAC systems.
@@
Against the aforementioned background, we are motivated to study an ISAC system aided by a fluid RIS (fRIS). In particular, fRIS, named after the fluid antenna system, implies that the elements of RIS are moveable. In this system, beams are formed at the fRIS to serve the communication users and illuminate the targets, where the sensing beampattern mismatch and communication decoding error are jointly minimized. This problem is solved by an alternating minimization (AM)-based algorithm, which incorporates augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite-relaxation (SDR), and majorization-minimization (MM) technique. The main contributions of this work are summarized as:
@@
  • As the movable elements can provide additional spatial DoFs compared to the conventional RIS, we propose a novel fRIS-aided ISAC system to improve the communication and sensing performance. Specifically, the ISAC transmit signal, fRIS beampattern for target illumination, communication symbol decoding and fRIS-related channels are characterized, and a joint sensing beampattern mismatch and communication decoding error minimization problem is formulated.
@@
  • To tackle the resultant optimization problem, we devise an algorithm based on AM. First, we reformulate the original high-order problem into a more manageable form. Then, the problem is divided into several subproblems and dedicated optimization approaches are established for each subproblem. Eventually, the convergence and the computational complexity of the proposed algorithm are analyzed.
@@
  • In the subproblem of optimizing the fRIS element positions, the optimization process is challenging since the changing RIS elements simultaneously affects the incident channel, the reflective channels. To address this, we leverage the characteristics of the fRIS and the structural channels to reduce the fourth-order terms of the element position to the linear and linear difference terms, so that the objective function of the subproblem can be simplified. This facilitates the optimization process.
  • Numerical results are demonstrated to validate the effectiveness of the proposed model. The results reveal a significant performance gain achieved by employing the fRIS in the ISAC system and the proposed scheme outperforms the benchmarks. Moreover, by employing the fRIS, the system performance loss caused by the reduction of the element number can be complemented, which alleviates the curse of dimensionality in the RIS optimization.
@@
This paper is organized as follows. Section II introduces the model architecture of the fRIS-aided ISAC system, while formulates the symbol error and beampattern mismatch minimization problem. In section III, the AM-based algorithm for the resultant problem and the analysis of the proposed algorithm are presented in details. Section IV demonstrates the numerical results. Finally, conclusions are drawn in Section V.
[…]
@@ II-A System Model
As depicted in Fig.1, we consider an ISAC system aided by a fRIS. Specifically, the ISAC base station (BS), equipped with a $M$-antenna uniform linear array (ULA), aims to simultaneously serve $K$ single-antenna communication users and sense $T$ point targets. Without loss of generalization, it is considered that the direct sensing and communication paths are blocked so that the fRIS is deployed to establish the virtual links and sensing paths. The fRIS has $N$ movable elements, where all the elements can move freely within a planar region $\mathcal{A}$ with the shape of $A\times A$. For $n$-th fRIS element, its position is denoted by its Cartesian coordinate $\mathbf{p}_{n}=\left[p_{x,n},p_{y,n}\right]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$. As a result, the positions of all the elements can be given compactly as $\mathbf{p}=\left[\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}\right]\in\mathcal{A}^{2\times N}$.
In this system, we denote the data symbols for the users as $\mathbf{s}_{c}\in\mathbb{C}^{K\times 1}$, which is precoded by a beamformer $\mathbf{W}\in\mathbb{C}^{M\times K}$ in the BS. This yields the precoded digital signal as
$\mathbf{x}=\mathbf{W}\mathbf{s}_{c}\in\mathbb{C}^{M\times 1}.$ (1)
The symbol $\mathbf{s}_{c}$ is considered statistically independent and auto-correlated, namely $\mathbb{E}[\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}]=\mathbf{I}_{K\times K}$. After converting the radio frequency domain, the signal $\mathbf{x}$ is emitted from the BS and propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS. The phase of the incident signal in the fRIS is altered adaptively by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\left\{\theta_{n},\forall n=1,\dots,N\right\}$ represents the phase shift coefficient for $n$-th RIS element. As a result, the reflective signal at the fRIS can be given as
@@
Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and propagate through the reflecting channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ to serve the communication users.
@@ II-A1 Sensing Model
As we know, the illumination power to the targets is direction related to the echo SNR, where more target information can obtained if the echo SNR is high. As a consequence, beams should be formed at the fRIS towards the directions of the targets to illuminate the targets. To this end, one can design the correlation matrix of the reflective signal $\mathbf{v}$ to manipulate the beampattern of the reflective signal. Specifically, the relationship between the beampattern and the correlation matrix of the reflective signal can be expressed as
@@
in which $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$ is the covariance matrix of signal $\mathbf{v}$. The symbol $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS, which is not only dependent on azimuth angle $\phi$ and elevate angle $\psi$, but also the positions of the movable elements $\mathbf{p}$. For conciseness, we leave the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in the subsection II-A3.
For beam design, a common metric is the beampattern similarity, where the mismatch between the practical beampattern $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and the ideal pattern $\mathcal{P}_{d}(\phi,\psi)$ are measured to evaluate the target sensing performance. Specifically, the beampattern mismatch can be written as
@@
in which $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote the grid number of the azimuth angles and elevate angles, respectively. When $\varepsilon_{r}$ is minimized, the practical beampattern aligns with the ideal beampattern, such that the reflective signal can effectively illuminate the targets.
@@ II-A2 Communication Model
Apart from illuminating the targets, the reflective signal $\mathbf{v}$ also propagates to the users for communication through the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$. The signal received at user $k$ can be given as
$y_{c,k}=\mathbf{h}_{rc,k}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+n_{c,k},~{}~{}\forall k=1,\cdots,K,$ (5)
where $n_{c,k}$ represents the additive Gaussian white noise at the receiver of user $k$ with zero mean and $\sigma_{0}^{2}$ variance. To estimate the transmitted symbols from the received signal, an estimation factor $\omega$ is introduced, i.e.,
$\hat{s}_{c,k}=\omega y_{c,k}.$ (6)
Given the estimated communication symbols, a direct performance indicator for evaluating the communication performance is the mean square error (MSE) between the estimated symbol and the transmitted symbol. It can be formulated as
$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$ (7a)
=$\overset{(a)}{=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (7b)
where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ incorporates all the reflective channels. The operation $(a)$ substitutes (5) and (6) to (7a) and leverages the statistical independence and the auto-correlation characteristics of $\mathbf{s}_{c}$.

### v3

@@ Abstract
A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communication (ISAC) system is proposed to enhance multi-target sensing and multi-user communication. Unlike the conventional RIS, the fRIS employs movable elements with adjustable positions, offering additional spatial degrees of freedom. In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and symbol estimation error. An algorithm based on alternating minimization is devised to handle the resultant non-convex problem, where the subproblems are solved via augmented Lagrangian method, quadratic programming, semidefinite relaxation, and majorization-minimization. A key challenge is that the element positions affect both incident and reflective channels, leading to the high-order composite objective functions. As a remedy, the high-order terms are transformed into linear and linear-difference forms by exploiting the structural characteristics of fRIS and the channels. Numerical results demonstrate the superiority of the proposed scheme over conventional RIS-aided ISAC and other benchmarks.
@@ I Introduction
In the era of sixth-generation wireless communications, integrated sensing and communications (ISAC) have emerged as a promising paradigm to alleviate resource scarcity [7]. In ISAC, the communication system and the sensing system aim to share the same spectrum [15], where various schemes have been developed, such as null-space projection [12] and transceiver design [30]. Furthermore, the works [36, 1] achieved collaborations of the two systems by sharing critical information. Additionally, given the hardware and signal processing similarities between sensing and communication, recent ISAC research focused on unified platforms capable of both functions [9]. For instance, the work [19] designed the transmit beamformer by combining radar waveforms and communication symbols, while the work [40] addressed wideband ISAC using a unified waveform design metric. In [25], a latency-aware resource allocation scheme was developed for multiple-input and multiple-output ISAC systems. In addition, recent works explored parameter estimation in ISAC systems. For example, the work [5] proposed a hybrid fusion framework, where the fused signal was separated for the estimation of the target directions.
Meanwhile, reconfigurable intelligent surfaces (RIS) have attracted great attention. Comprised of many reflective elements, RIS is capable of inducing phase shifts on incident signals and dynamically reconfiguring wireless channels [24]. This capability enables substantial communication sum-rate gains in multi-user scenarios [35], and facilitates suppression of undesired signals such as eavesdropping [8] and jamming [33]. Besides, the work [26] showed that RIS can construct favorable channels to support devices with ultra-reliable and low-latency communications. In addition, RIS can be applied to various areas, such as radar detection [29], near-field communications [44], and wireless power transfer and communication [45].
To harness the combined benefits of ISAC and RIS, extensive studies have been conducted on RIS-aided ISAC systems [18]. In [10], the authors maximized the echo signal-to-noise ratio (SNR) by jointly designing the transmitter and RIS beamformer in a single-user and single-target scenario. This framework was extended in [17] to support multi-user and multi-target systems. Furthermore, the work [2] addressed the issue of high dynamic range caused by strong interference in RIS-aided ISAC systems, where a joint active and passive beamforming was designed to reduce the dynamic range and mitigate the interference. To reduce energy consumption, energy efficiency optimization was explored in RIS-aided ISAC systems [55, 48]. Additionally, different applications were investigated, including simultaneously transmitting and reflecting surfaces [39], unsupervised learning [47], and extended reality [22].
@@
In the above works, there remains potential to improve spatial degrees of freedom (DoFs) and reduce system complexity. Particularly, the conventional arrays have fixed antennas after assembly, which restricts the spatial diversity [23]. Furthermore, achieving high performance often requires a large number of elements, leading to increased manufacturing costs and optimization burdens. To overcome these issues, the fluid antenna (FA), also known as movable antenna, has emerged as a candidate. Enabled by fluidic metamaterials, FA can dynamically adjust the antenna positions, offering additional spatial DoFs [41]. Moreover, the improved performance enables comparable gains with fewer elements, thereby reducing the computational load of large-scale optimization. The potential of FA was extensively explored, such as point-to-point capacity [23], communication sum-rate [27], and ISAC performance improvement [56, 58, 51].
To adaptively control the propagation environment, some efforts incorporated RIS in FA systems [31, 46]. However, they still rely on fixed RIS elements. Considering that RIS has more elements, fluid RIS (fRIS) or movable-element RIS was studied [32, 43, 6, 57, 53, 54, 20] to exploit more DoFs, where the elements in RIS are movable. Compared to the RIS-aided FAS system, fRIS has several advantages. First, fRIS can bring more spatial DoFs, as the RIS typically contains more elements than the BS antennas. Second, RIS elements are passive and free from the RF chain constraints. Third, from a deployment perspective, fRISs are typically more ubiquitous than FAS-BSs in large-scale networks, effectively improving system performance.
In the fRIS systems, several works [32, 43, 6] explored pixel-based hardware architectures that emulate virtual element movement through densely distributed port switching. In contrast, other works [57, 53, 54, 20] employed flexible cables, motors, and sliding tracks to reposition the elements for communication enhancement. While these implementations inevitably add hardware complexity, the fRIS can achieve improved performance with fewer elements due to the additional DoFs in a new dimension [^1] The fRIS introduces new spatial DoFs by allowing controlled displacement of each reflecting element as compared to the traditional RIS. These position-based DoFs are fundamentally distinct from phase-shift control, as they reshape the propagation geometry, offering a complementary optimization domain. . Additionally, enabling technologies are actively studied and continuously refined, making the hardware complexity manageable. This offers a new direction for RIS evolution. However, these works concerning the fRIS mainly focus on basic communication settings and lack ISAC modeling and optimization, motivating further study in this direction.
@@
Against the above background, we investigate a fRIS-aided ISAC system, where the fRIS forms beams to simultaneously serve communication users and illuminate sensing targets. A joint optimization problem is formulated to minimize both sensing beampattern mismatch and communication decoding error. Then, the resultant problem is addressed using the alternating minimization (AM)-based algorithm incorporating augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite relaxation (SDR), and majorization-minimization (MM). The main contributions are summarized as follows:
@@
  • Leveraging the additional spatial DoFs introduced by movable elements, we propose a novel fRIS-aided ISAC system to enhance sensing and communication. Specifically, we characterize the transmit signal, fRIS beampattern, symbol decoding, and fRIS-related channels. Besides, we formulate a joint minimization problem for sensing mismatch and decoding error.
@@
  • To solve the resultant problem, we devise an AM-based algorithm. First, the original high-order formulation is transformed into a more tractable form. Then, the problem is decomposed into subproblems, which are tackled using dedicated optimization techniques. Furthermore, we analyze the convergence and computational complexity of the proposed algorithm.
@@
  • The subproblem of optimizing fRIS element positions is particularly challenging, as their movement affects both incident and reflective channels. To address this issue, we exploit the structural properties of fRIS and channels to reduce the fourth-order terms to linear and linear-difference forms, simplifying the objective and facilitating the optimization.
  • Numerical results validate the effectiveness of the proposed model, showing notable performance gains over existing benchmarks. Moreover, fRIS compensates for performance degradation caused by reducing the number of RIS elements, alleviating the high-dimensional complexity in RIS optimization.
@@
This paper is organized as follows. Section II presents the architecture of the fRIS-aided ISAC system and formulates the joint minimization problem. Section III details the proposed AM-based algorithm, while analyzes its convergence and complexity. Numerical results are shown in Section IV, and conclusions are drawn in Section V.
Notation: Bold lowercase and uppercase letters denote vectors and matrices, respectively. In addition, $(\cdot)^{\mathrm{H}}$, $(\cdot)^{\mathrm{T}}$, and $(\cdot)^{*}$ signify conjugate transpose, transpose, and conjugate, respectively. $\mathrm{tr}(\cdot)$ is the trace of a matrix. $\mathrm{diag}(\cdot)$ extracts diagonal entries into a vector, while $\mathrm{Diag}(\cdot)$ creates a diagonal matrix from a vector. $\mathrm{vec}(\mathbf{A})$ vectorizes matrix $\mathbf{A}$. $|\cdot|$ and $\|\cdot\|$ denote absolute value and norm, respectively. $\mathfrak{R}[x]$ and $\angle x$ are the real part and phase of a complex scalar $x$, respectively.
@@ II-A System Model
As illustrated in Fig. 1, we consider an ISAC system assisted by a fRIS. The ISAC base station (BS), equipped with a uniform linear array (ULA) of MM antennas, simultaneously serves $K$ single-antenna communication users and senses $T$ point targets. Without loss of generality, we assume that direct links for both sensing and communication are blocked and thus a fRIS is deployed to establish virtual communication and sensing paths. The fRIS is comprised of $N$ movable elements, where each element can reposition freely within a planar region $\mathcal{A}$ of size $A\times A$. The position of the $n$-th element is denoted as $\mathbf{p}_{n}=[p_{x,n},p_{y,n}]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$, where all the element positions are given by $\mathbf{p}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{N}]\in\mathcal{A}^{2\times N}$.
In this system, the signal $\mathbf{x}\in\mathbb{C}^{M}$ is transmitted from the BS, which acts as dual-functional waveform for both sensing and communication. The transmit signal propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS. In the fRIS, the phase of the incident signal $\mathbf{G}\mathbf{x}$ is altered by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\theta_{n}$ represents the phase shift coefficient for the $n$-th fRIS element. As a result, the reflective signal at the fRIS is given by
@@
Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and serve the communication users via the reflecting channel $\mathbf{H}_{rc}$, where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ represents the overall downlink reflective channel while $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ is the reflective channel from the fRIS to the $k$-th user.
@@ II-A1 Sensing Model
In the fRIS, beams should be formed towards the directions of the targets for target illuminations. To this end, we can design $\mathbf{v}$ to manipulate the beampattern of the reflective signal. Specifically, the beampattern of $\mathbf{v}$ is expressed as
@@
which indicates the illumination power of the fRIS at the azimuth angle $\phi$ and the elevation angle $\psi$. In (2), we define $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$, while the symbol of $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS. For conciseness, we present the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in Subsection II-A3.
For beam design, a common metric is the beampattern similarity, which evaluates the mismatch between $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and an ideal pattern $\mathcal{P}_{d}(\phi,\psi)$. The mismatch is quantified as
@@
where $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote azimuth and elevation angle sample numbers, respectively. Minimizing $\varepsilon_{r}$ encourages $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ to closely match $\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})$, enhancing the target illumination performance.
@@ II-A2 Communication Model
After $\mathbf{v}$ propagates to the communication users, the overall received signal $\mathbf{y}_{c}\in\mathbb{C}^{K}$ at the users is given by
$\mathbf{y}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+\mathbf{n}_{c},$ (4)
where $\mathbf{n}_{c}\in\mathbb{C}^{K}$ represents the additive Gaussian white noise matrix, whose entries are independent and identically distributed with zero mean and $\sigma_{0}^{2}$ variance. We denote $y_{c,k}$ as the received signal of the $k$-th user, which is the $k$-th entry of $\mathbf{y}_{c}$. Given a desired constellation symbol ${{s}}_{c,k}$ for the $k$-th user, an estimation factor $\omega$ is introduced to recover ${{s}}_{c,k}$ from $y_{c,k}$, which is given by
$\hat{s}_{c,k}=\omega y_{c,k},$ (5)
where $\hat{s}_{c,k}$ is the estimated symbol. To evaluate the communication performance, a direct metric is the mean square error (MSE) between the estimated symbol and the transmitted symbol:
$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$
${=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (6)
where $\mathbf{s}_{c}\in\mathbb{C}^{K}$ and $\hat{\mathbf{s}}_{c}\in\mathbb{C}^{K}$ incorporate all the ${{s}}_{c,k}$ and $\hat{s}_{c,k}$, respectively. In (II-A2), the first term can be viewed as the multi-user interference, while the second term is the impact of the noise.